OpenStudy (mysangels):

Please help! A square has an area of 500 square units. Find the length of a side of the square as a radical expression in simplest form.

OpenStudy (theopenstudyowl):

Hey! Let me just review some things first. Been a while since I have done geometry.

OpenStudy (mysangels):

This problem was in the unit for square roots and such.

OpenStudy (anonymous):

take the square root that is all

OpenStudy (theopenstudyowl):

thats it?

OpenStudy (theopenstudyowl):

lol

OpenStudy (mysangels):

So, just find the square root of 500?

OpenStudy (anonymous):

well they want "radical form" so you cannot write \[\sqrt{500}\] or a decimal

OpenStudy (mysangels):

You don't need to have a 4 outside of the radical sign?

OpenStudy (anonymous):

since \[500=5\times 100\] you know \[\sqrt{500}=\sqrt{100\times 5}=\sqrt{100}\sqrt5=10\sqrt5\]

OpenStudy (theopenstudyowl):

10 radical 5 is right!

OpenStudy (mysangels):

Thank you so much @satellite73 and @theopenstudyowl ! Can you guys help me with one more?

OpenStudy (theopenstudyowl):

I can...

OpenStudy (mysangels):

Thank you!

OpenStudy (theopenstudyowl):

I can help you with as many as you want!

OpenStudy (mysangels):

You stand on a bridge and drop a pebble into the water below from a height of 75 feet. You can use the expression \[\sqrt{\left(\begin{matrix}75 \\ 16\end{matrix}\right)}\] to find the time in seconds that it takes the pebble to hit the water. Write the expression in simplest form. Then approximate the value of the expression to the nearest second.

OpenStudy (theopenstudyowl):

what class is this?

OpenStudy (mysangels):

pre-algebra

OpenStudy (theopenstudyowl):

haha, its just I dont remember ever seeing a notation like that one, thats all... one sec.

OpenStudy (theopenstudyowl):

is that a fraction?

OpenStudy (mysangels):

yes

OpenStudy (theopenstudyowl):

oh ok, that makes sense!

OpenStudy (mysangels):

yes

OpenStudy (mysangels):

Can you fix the equations? I lost you.

OpenStudy (theopenstudyowl):

yes, one sec

OpenStudy (mysangels):

Thank you!

OpenStudy (theopenstudyowl):

\[\sqrt{75/16}=\sqrt{75}/\sqrt{16}=\sqrt{75}/4=(\sqrt{25}\times \sqrt{3}/4=5\sqrt{3}/4\]

OpenStudy (mysangels):

Would it be 25 x 3/4 because 25x3= 75, right?

OpenStudy (theopenstudyowl):

yes!

OpenStudy (theopenstudyowl):

then the square root of 25 is 5

OpenStudy (mysangels):

Ok. Thank you!

OpenStudy (mysangels):

How would you find the value of the expression to the nearest second?

OpenStudy (theopenstudyowl):

plug the 5sqrt3/4 into a calculator and round to the nearest whole number..

OpenStudy (mysangels):

Could you work out the problem, though?

OpenStudy (theopenstudyowl):

2 seconds!

OpenStudy (theopenstudyowl):

the lowest possible form is \[5\sqrt{3}/4\]

OpenStudy (mysangels):

Is that the approximate second?

OpenStudy (theopenstudyowl):

i believe so!

OpenStudy (mysangels):

Ok. Thank you for all of your help!

OpenStudy (theopenstudyowl):

yah np

OpenStudy (mysangels):

Bye!

OpenStudy (theopenstudyowl):

see yah